# TokenWeave: Efficient Compute-Communication Overlap for Distributed LLM Inference

### v1

@@ 3.1.2. Overlapped Execution
The operations of these two batches can now be overlapped as shown in Figure 8(b). As shown, when the AllReduce of the first batch is being processed, we compute the attention of the second batch. The FFN of the first batch is then overlapped with the AllReduce of the second batch and so on. We implement this overlapped execution via CUDA streams. A communication stream handles all the communication operations, while a compute stream runs the compute operations. Lightweight synchronization using CudaEventSynchronize is performed between the compute and communication stream to handle data dependencies.
@@ 3.2. RMSNorm Reordering
As described in Section 2.1, distributed inference via tensor-parallelism involves an AllReduce operation after Attention and FFN computations. In transformer models, the RMSNorm operation typically follows these layers and is computed independently by each GPU. This independent computation results in redundancy since all GPUs have identical token embeddings post-AllReduce.
To address this inefficiency, TokenWeave strategically reorders RMSNorm operations within the AllReduce process. Specifically, the AllReduce operation can be decomposed into ReduceScatter and AllGather operations. At the completion of the ReduceScatter step, each GPU possesses the complete and final state of $\frac{1}{N}$ of the tensor, where $N$ represents the total GPUs in the distributed replica. Consequently, each GPU can independently perform RMSNorm on its dedicated $\frac{1}{N}$ portion without redundancy. Since RMSNorm is a token level operation, we only need to ensure that ReduceScatter splits the tensor only at token boundaries. This ensures that each GPU has the full token embeddings. The subsequent AllGather operation then distributes these post-RMSNorm values to all GPUs.
Through this reordering, TokenWeave reduces RMSNorm computation by a factor of $N$ compared to traditional implementations, eliminating unnecessary redundancy. However, as shown in Figure 4, such a simple reordering can actually result in performance loss as the cost of dividing AllReduce into separate ReduceScatter and AllGather cancels the RMSNorm computation gains. We address this inefficiency through our fused kernel in the next section.
@@
Listing 1: Implementation of Fused AllReduce+RMSNorm kernel. Current implementation only supports BFloat16. For brevity, some parts of the code have been omitted. The fused kernel is built on top of the PyTorch MultiMem AllReduce code (PyTorch Team, [n. d.]) and the RMSNorm kernel from vLLM (vLLM Contributors, [n. d.]).
3.3. Fused AllReduce-RMSNorm Implementation
@@
The reduction in memory accesses, combined with the elimination of redundant computations, makes our fused AllReduce-RMSNorm kernel very efficient. As shown in Figure 4, the fused kernel provides up to 40% gains over the current approach of separate AllReduce + RMSNorm computation across all sequence lengths.
Also, the reduced compute and memory bandwidth requirements, allows this fused kernel to be executed with very few SMs (only 2-8 in our experiments) without incurring much overhead — as shown in Figure 10 the kernel execution time does not improve much after 8 SMs and for larger sequences, even 4 SMs are enough. This allows us to overlap the full AllReduce-RMSNorm operation of one split-batch with the compute of the other split, enabling overlap of not just the communication but also the memory-bound RMSNorm. Since this computation is overlapped with the other split-batch, even if using fewer SMs (e.g. 2 instead of 8) results in some overheads, it will be hidden as long as its execution time does not exceed the execution time of the overlapped compute kernel.
Figure 10. The fused AllReduce-RMSNorm kernel performs optimally with very few SMs. Shown is the performance of AllReduce multimem kernel under varying number of SMs for different sequence lengths (hidden size 8192, bf16). 8 SMs is close to optimal in most cases.
@@ 4.1. Experimental Setup
Implementation Details: We implemented TokenWeave over the vLLM (Kwon et al., 2023) framework, a state-of-the-art serving system for LLM inference. Specifically, we build on the recent and highly performant V1 engine (vLLM Team, [n. d.]) available in vLLM==v0.8.5.dev106. We leverage PyTorch==v2.6.0, which includes support for SymmetricMemory and implemented the communication collectives using triton-3.2.0 and PyTorch CUDA extensions built on top of symmetric memory recipes recently released by PyTorch (Wang et al., 2025).
@@
Environment: All our experiments are performed on an 8xH100 NVIDIA DGX system with NV-SHARP support. The TP-4 experiments use the same system but use four of the eight GPUs. We also use fixed TDP settings to ensure stable performance measurements (NVIDIA, 2020). For microbenchmarks, we perform warm-up, flush the L2 cache and use cudaEvents for timing measurements. For end-to-end (e2e) results, we perform warm-up and then use time.perf_counter() for timing. We disable prefix caching to ensure consistent and fair comparisons. For end-to-end evaluations, we ignore the detokenization time (time to convert output tokens back to text performed in the CPU).
@@ 4.2. Experimental Results
Figure 11. TokenWeave Latency Gains. Shown are the execution times of a prefill request with varying sequence length for different models on (a) 8xH100s and (b) 4xH100s. In almost all cases, TokenWeave is close to or better than the theoretical vllm-nocomm baseline with zero communication overheads, showing that TokenWeave not only recovers all the communication overheads, but provides additional gains due to RMSNorm fusion.
4.2.1. TokenWeave Latency Benefits
First, we evaluate latency for a single iteration of the forward pass. For this experiment, we use prefill-only batches with varying sequence lengths as shown in Figure 11. We evaluate five different settings, corresponding to three models running on 8xH100s and 4xH100s, except for Mixtral-8x22B that does not fit on 4xH100s. As shown, TokenWeave shows consistent gains between $18-29\%$ for the dense models in the 8xH100s case, with significant gains even at small sequence lengths. For the MoE Mixtral model, the gains are generally smaller but the communication overhead for Mixtral is also lower (Figure 1). Further, unlike the dense models, there is a net overhead for TokenWeave in Mixtral at smaller sequence lengths of 1K and 2K. This is because, in MoE, the tokens for MLP computation get split across the 8 experts, making the FFN computations memory bound for smaller requests. As a result, even TokenWeave’s coarse splitting into two sub-tasks can result in non-trivial overheads that cannot be compensated by saving due to computation overlap. Finally, for the 4xH100s experiments, TokenWeave again gets substantial gains consistently over the range of sequence length starting from 1K. The gains are lower than 8xH100s, as the communication overhead itself is less with fewer GPUs.
(a) 8xH100
(b) 4xH100
Figure 12. TokenWeave throughput gains for end-to-end workload traces. Shown is the measured throughput under fixed (input, output) length traces as well as the ShareGPT trace for different models on (a) 8xH100s and (b) 4xH100s.
@@ 4.2.1. TokenWeave Latency Benefits
Figure 13. TokenWeave throughput gains for end-to-end workload traces under chunk size variation. Shown is the measured throughput under fixed (input, output) length traces as well as the ShareGPT trace for Llama-3.3-70B on 8xH100 DGX. The chunk-size is varied from 1024 to 8192.
@@ 4.2.2. TokenWeave Throughput Gains
We now evaluate the throughput performance of TokenWeave for various workload traces. As is standard practice, we use chunked-prefills (Agrawal et al., 2024) with hybrid prefill-decode batching to ensure consistent latency guarantees while maximizing throughput. Chunked-prefill ensures that time-between-tokens (TBT) latency is bounded, resulting in smooth response to the user. Chunked-prefills is turned on by default in the vLLM framework. For the first experiment, we use a chunk size of 2K (vLLM default) for the dense models and 4K for Mixtral. We apply TokenWeave to hybrid batches with 1K or more tokens (4K for Mixtral) and use non-overlapped communication using our fused kernel for the smaller decode-only batches.
Figure 12(b) shows the performance evaluation with chunked-prefills on workloads with both fixed input and output length requests, as well as ShareGPT (Wang et al., 2023) which is a real-world workload trace with varying input/output requests lengths based on user conversations. As shown, in the 8xH100s case, TokenWeave continues to show substantial throughput gains of $\sim 20\%$ for the dense models and recovers most of the communication overheads.
The chunk size used in chunked-prefill enables a trade-off between TBT latency and throughput, with smaller chunk sizes enabling lower TBT values but also resulting in lower throughput. In the second experiment, we vary the chunk size and evaluate TokenWeave’s throughput. Figure 13 uses the same workload as before but with varying chunked-prefill sizes for the Llama-3.3-70B model. The figure shows that there is consistent performance gains of 15-26% with TokenWeave over a range of chunk sizes.
@@ 4.2.3. Comparison with TileLink
For this evaluation, we use batch size 1 and vary the sequence length of the single request. Figure 14 shows the time taken for a single layer computation. As shown, TokenWeave reduces the computation latency by $20\%$ even at a small request length of 1K tokens, while state-of-the-art TileLink actually ends up with a net overhead at small sequence lengths. TileLink results in performance improvements only for sequence lengths $>4K$, but the gains saturate to around $21\%$, while TokenWeave provides much higher gains up to $38\%$. Note that end-to-end latency gains in Section 4.2.1 are lower than the single-layer latency gains here due to non-layer overheads like embedding, sampling, etc.
@@ 4.2.4. Comparison with NanoFlow
While the NanoFlow paper shows that their serving stack is more efficient than an earlier version of vLLM on A100 GPUs, we find our vLLM V1 baseline throughput numbers on H100 GPUs are better than NanoFlow’s serving stack, perhaps because vLLM V1 has significantly improved its performance recently. Thus, instead of performing an absolute throughput comparison, we resort to a relative performance comparison as described next.
We evaluate the adapted NanoFlow implementation under two conditions: with NanoFlow enabled and disabled within its own serving stack. Figure 15 shows the performance evaluation with chunked-prefills (chunk size 2K) on workloads with various fixed input and output length requests. As shown, NanoFlow provides only modest performance gains in the range of 5-8%, closely matching the 7% communication gains reported in the NanoFlow paper. In contrast, TokenWeave demonstrates significantly larger gains ($\sim 20\%$) across a broad range (see Figure 12(b)).
@@ 5. Conclusion
We find the communication cost for large models served over multiple GPUs is as high as 20% today despite hardware support such as high speed NVLink and NVSHARP. Furthermore, we identify that RMSNorm also results in a significant overhead of 5–9%. TokenWeave addresses these issues by splitting the model input into two approximately equal batches and overlapping the compute of one batch with a novel fused AllReduce-RMSNorm communication and normalization kernel of the other batch. Through extensive experimental evaluations using multiple models running on 4 H100 and 8 H100 DGX GPUs, we show that TokenWeave provides up to 29% latency gains and up to 26% throughput gains compared to an optimized baseline in a variety of settings and workloads.

### v2

@@ 3.1.2. Overlapped Execution
The operations of these two batches can now be overlapped as shown in Figure 8(b). As shown, when the AllReduce of the first batch is being processed, we compute the attention of the second batch. The FFN of the first batch is then overlapped with the AllReduce of the second batch and so on. We implement this overlapped execution via CUDA streams. A communication stream handles all the communication operations, while a compute stream runs the compute operations. Lightweight synchronization using CudaStreamSynchronize is performed between the compute and communication stream to handle data dependencies.
@@ 3.2. RMSNorm Reordering
As described in Section 2.1, distributed inference via tensor-parallelism involves an AllReduce operation after Attention and FFN computations. In transformer models, the Residual Addition and RMSNorm operation typically follows these layers and is computed independently by each GPU. This independent computation results in redundancy since all GPUs have identical token embeddings post-AllReduce.
To address this inefficiency, TokenWeave strategically reorders RMSNorm operations within the AllReduce process. Specifically, the AllReduce operation can be decomposed into ReduceScatter and AllGather operations. At the completion of the ReduceScatter step, each GPU possesses the complete and final state of $\frac{1}{N}$ of the tensor, where $N$ denotes the total number of GPUs in the distributed replica. Consequently, each GPU can independently perform RMSNorm on its dedicated $\frac{1}{N}$ portion without redundancy. Since RMSNorm is a token level operation, we only need to ensure that ReduceScatter splits the tensor only at token boundaries. This ensures that each GPU has the full token embeddings. The subsequent AllGather operation then distributes these post RMSNorm values to all GPUs.
Through this reordering, TokenWeave reduces the RMSNorm computation by a factor of $N$ compared to traditional implementations, thereby eliminating unnecessary redundancy. However, as shown in Figure 4, such a simple reordering can actually result in performance loss as the cost of dividing AllReduce into separate ReduceScatter and AllGather cancels the RMSNorm computation gains. We address this inefficiency through our fused kernel in the next section.
@@
Listing 1: Implementation of Fused AllReduce–RMSNorm kernel. Current implementation only supports BFloat16. For brevity, some parts of the code have been omitted. The fused kernel is built on top of the PyTorch MultiMem AllReduce code (PyTorch Team, [n. d.]) and the RMSNorm kernel from vLLM (vLLM Contributors, [n. d.]).
3.3. Fused AllReduce–RMSNorm Implementation
@@
The reduction in memory accesses, combined with the elimination of redundant computations, makes our fused AllReduce–RMSNorm kernel very efficient. As shown in Figure 4, the fused kernel provides up to 1.40×1.40$\times$1.40 × improvement over the current approach of separate AllReduce + RMSNorm computation across all sequence lengths.
Also, the reduced compute and memory bandwidth requirements, allows this fused kernel to be executed with very few SMs (only $2-8$ in our experiments) without incurring much overhead — as shown in Figure 10 the kernel execution time does not improve much after $8$ SMs and for larger sequences, even $4$ SMs are enough. This allows us to overlap the full AllReduce–RMSNorm operation of one split-batch with the compute of the other split, enabling overlap of not just the communication but also the memory-bound RMSNorm. Since this computation is overlapped with the other split-batch, even if using fewer SMs (e.g. $2$ instead of $8$) results in some overheads, it will be hidden as long as its execution time does not exceed the execution time of the overlapped compute kernel.
Figure 10. The fused AllReduce–RMSNorm kernel performs optimally with very few SMs. Shown is the performance of AllReduce multimem kernel under varying number of SMs for different sequence lengths (hidden size $8192$, $bf16$). $8$ SMs is close to optimal in most cases.
@@ 4.1. Experimental Setup
Implementation Details: We implemented TokenWeave over the vLLM (Kwon et al., 2023) framework, a state-of-the-art serving system for LLM inference. Specifically, we build on the recent and highly performant V1 engine (vLLM Team, [n. d.]) available in vLLM==v0.8.5. We leverage PyTorch==v2.6.0, which includes support for SymmetricMemory and implemented the communication collectives using triton-3.2.0 and PyTorch CUDA extensions built on top of symmetric memory recipes recently released by PyTorch (Wang et al., 2025).
@@
Environment: All our experiments are performed on an 8×8$\times$8 ×H100 NVIDIA DGX system with NV-SHARP support. The $TP-4$ experiments use the same system but use four of the eight GPUs. We also use fixed TDP settings to ensure stable performance measurements (NVIDIA, 2020). For microbenchmarks, we perform warm-up, flush the L2 cache and use cudaEvents for timing measurements. For end-to-end (e2e) results, we perform warm-up and then use time.perf_counter() for timing. We disable prefix caching to ensure consistent and fair comparisons. For end-to-end evaluations, we ignore the detokenization time (time to convert output tokens back to text performed in the CPU).
@@ 4.2. Experimental Results
(a) 8×8$\times$8 ×H100
(b) 4×4$\times$4 ×H100
Figure 11. TokenWeave Latency Gains. Shown are the execution times of a prefill request with varying sequence length for different models on (a) 8×8$\times$8 ×H100 and (b) 4×4$\times$4 ×H100. In almost all cases, TokenWeave is close to or better than the theoretical vllm-nocomm baseline with zero communication overheads, showing that TokenWeave not only recovers all the communication overheads, but provides additional gains due to RMSNorm fusion.
4.2.1. TokenWeave Latency Benefits
First, we evaluate latency for a single iteration of the forward pass. For this experiment, we use prefill-only batches with varying sequence lengths as shown in Figure 11. We evaluate five different settings, corresponding to three models running on 8×8$\times$8 ×H100 and 4×4$\times$4 ×H100, except for Mixtral-8x22B that does not fit on 4×4$\times$4 ×H100. As shown, TokenWeave consistently achieves between 1.18×1.18$\times$1.18 × and 1.29×1.29$\times$1.29 × improvement for dense models on 8×8$\times$8 ×H100, with notable gains even at shorter sequence lengths. For the MoE Mixtral model, the gains are generally smaller but the communication overhead for Mixtral is also lower (Figure 1). Further, unlike the dense models, there is a net overhead for TokenWeave in Mixtral at smaller sequence lengths of $1K$ and $2K$. This is because, in MoE, the tokens for MLP computation get split across the $8$ experts, making the FFN computations memory bound for smaller requests. As a result, even TokenWeave’s coarse splitting into two sub-tasks can result in non-trivial overheads that cannot be compensated by saving due to computation overlap. Finally, for the 4×4$\times$4 ×H100 experiments, TokenWeave again gets substantial gains consistently over the range of sequence length starting from $1K$. The gains are lower than 8×8$\times$8 ×H100, as the communication overhead itself is less with fewer GPUs.
@@ 4.2.1. TokenWeave Latency Benefits
Figure 12. TokenWeave throughput gains for end-to-end workload traces. Shown is the measured throughput under fixed (input, output) length traces as well as the ShareGPT trace for different models on (a) 8×8$\times$8 ×H100 and (b) 4×4$\times$4 ×H100.
@@
Figure 13. TokenWeave throughput gains for end-to-end workload traces under chunk size variation. Shown is the measured throughput under fixed (input, output) length traces as well as the ShareGPT trace for Llama-3.3-70B on 8×8$\times$8 ×H100 DGX. The chunk-size is varied from $1024$ to $8192$.
@@ 4.2.2. TokenWeave Throughput Gains
We now evaluate the throughput performance of TokenWeave for various workload traces. As is standard practice, we use chunked-prefills (Agrawal et al., 2024) with hybrid prefill-decode batching to ensure consistent latency guarantees while maximizing throughput. Chunked-prefill ensures that time-between-tokens (TBT) latency is bounded, resulting in smooth response to the user. Chunked-prefills is turned on by default in the vLLM-V1 framework. For the first experiment, we use a chunk size of $2K$ (vLLM default) for the dense models and $4K$ for Mixtral. We apply TokenWeave to hybrid batches with $1K$ or more tokens ($4K$ for Mixtral) and use non-overlapped communication using our fused kernel for the smaller decode-only batches.
Figure 12(b) shows the performance evaluation with chunked-prefills on workloads with both fixed input and output length requests, as well as ShareGPT (Wang et al., 2023) which is a real-world workload trace with varying input/output requests lengths based on user conversations. As shown, in the 8×8$\times$8 ×H100 case, TokenWeave continues to show substantial throughput gains of approximately 1.20×1.20$\times$1.20 × for dense models, effectively recovering most of the communication overhead.
The chunk size used in chunked-prefill enables a trade-off between TBT latency and throughput, with smaller chunk sizes enabling lower TBT values but also resulting in lower throughput. In the second experiment, we vary the chunk size and evaluate TokenWeave’s throughput. Figure 13 uses the same workload as before but with varying chunked-prefill sizes for the Llama-3.3-70B model. The figure shows consistent performance improvements of 1.15×1.15$\times$1.15 × to 1.26×1.26$\times$1.26 × with TokenWeave across a range of chunk sizes.
@@ 4.2.3. Comparison with TileLink
For this evaluation, we use batch size $1$ and vary the sequence length of the single request. Figure 14 shows the time taken for a single layer computation. As shown, TokenWeave achieves a 1.20×1.20$\times$1.20 × latency speedup even at a small request length of $1K$ tokens, while state-of-the-art TileLink actually ends up with a net overhead at small sequence lengths. TileLink results in performance improvements only for sequence lengths $>4K$, but the improvements saturate at around 1.21×1.21$\times$1.21 ×, while TokenWeave provides significantly higher improvements of up to 1.38×1.38$\times$1.38 ×. Note that the end-to-end latency gains reported in Section 4.2.1 are lower than the single-layer gains shown here due to additional non-layer overheads such as embedding, sampling, and other runtime components.
@@ 4.2.4. Comparison with NanoFlow
While the NanoFlow paper shows that their serving stack is more efficient than an earlier version of vLLM on A100 GPUs, we find our vLLM-V1 baseline throughput numbers on H100 GPUs are better than NanoFlow’s serving stack, perhaps because vLLM-V1 has significantly improved its performance recently. Thus, instead of performing an absolute throughput comparison, we resort to a relative performance comparison as described next.
We evaluate the adapted NanoFlow implementation under two conditions: with NanoFlow enabled and disabled within its own serving stack. Figure 15 shows the performance evaluation with chunked-prefills (chunk size $2K$) on workloads with various fixed input and output length requests. As shown, NanoFlow provides only modest performance gains in the range of of 1.05×1.05$\times$1.05 × to 1.08×1.08$\times$1.08 ×, closely matching the 1.07×1.07$\times$1.07 × communication improvement reported in the NanoFlow paper. In contrast, TokenWeave demonstrates significantly larger improvements of approximately 1.20×1.20$\times$1.20 × across a broad range of scenarios (see Figure 12(b)).
@@
4.2.5. AllReduce–RMSNorm Fusion Ablation
(a) 8×8$\times$8 ×H100
(b) 4×4$\times$4 ×H100
Figure 16. TokenWeave Ablation. Shown are the execution times of a prefill request with varying sequence lengths for different models on (a) 8×8$\times$8 ×H100 and (b) 4×4$\times$4 ×H100. TokenWeave-fuseonly provides gains due to the elimination of redundancy in RMSNorm computation and intermediate memory accesses; while TokenWeave provides additional gains from the compute-communication overlap.
To quantify the performance contributions of our optimizations, we perform an ablation study comparing three configurations: vllm-Multimem (baseline), TokenWeave-fuseonly (fused kernel only without the batch splitting and compute-communication overlap), and full TokenWeave (fused kernel with the compute-communication overlap). Figure 16 reports the latency of a single prefill request across varying sequence lengths, evaluated over multiple models under both 8$\times$×H100 and 4$\times$×H100 settings.
TokenWeave-fuseonly demonstrates significant gains across all models and hardware configurations. This is due to the eliminiation of redundant RMSNorm computations and intermediate memory accesses, as well as the reduced kernel launch overhead. For example, in the Llama-3.3-70B-Instruct (8$\times$×H100) setting, we observe up to 1.09×1.09$\times$1.09 × speedup at $8K$ tokens compared to vllm-Multimem. Similar trend is seen for Qwen2.5-72B-Instruct and Mixtral-8x22B, where TokenWeave-fuseonly provides consistent gains in the range of 1.041.09×1.04-1.09$\times$1.04 - 1.09 ×. The 4$\times$×H100 configuration shows consistent gains as well, although lower than the 8$\times$×H100 setting. This is because the redundancy in RMSNorm computation decreases (which TokenWeave-fuseonly eliminates) as the TP dimension decreases.
Additionally, TokenWeave provides substantial gains over TokenWeave-fuseonly due to the overlapped communication. We see consistent gains across models and hardware configurations. The gains in 8$\times$×H100 are higher than in 4$\times$×H100 as the communication overhead itself reduces with fewer GPUs. In the MoE Mixtral model, the gains are relatively smaller and TokenWeave has a net overhead at smaller sequence lengths of 1K and 2K due to the increased splitting overhead from the memory bound FFN operations in MoE models at smaller tokens, as discussed in section 4.2.1.
@@ 5. Conclusion
We find that the communication cost for large models served over multiple GPUs is as high as $20\%$ today despite hardware support such as high-speed NVLink and NVSHARP. Furthermore, we identify that RMSNorm also results in a significant overhead of $5-9\%$. TokenWeave addresses these issues by splitting the model input into two approximately equal batches and overlapping the compute of one batch with a novel fused AllReduce–RMSNorm communication and normalization kernel of the other batch. Through extensive experimental evaluations using multiple models running on 4×4$\times$4 ×H100 and 8×8$\times$8 ×H100 DGX GPUs, we show that TokenWeave achieves up to 1.29×1.29$\times$1.29 × faster latency and 1.26×1.26$\times$1.26 × higher throughput compared to an optimized baseline in a variety of settings and workloads.